Problem 21.11. Prove the following corollary of Theorem 21.13 above.
Theorem 21.13. Let A, B,C, and D be nonempty sets with AC and Bn D. Then

Hello, can you please solve 21.11, using the Theorem 21.13? Thank you.
Problem 21.11. Prove the following corollary of Theorem 21.13 above.
Theorem 21.13. Let A, B,C, and D be nonempty sets with AC and Bn D. Then
Answer #1

Defination : we say that two set A \approx B if there exist a bijective map f:A→B .

Given , A \approx C , B \approx D , so there exist two map f:A→C and g: B\rightarrow D which are bijective .

Now , h: A\times B\rightarrow C\times D be defined by ,

h ( x , y ) = ( f(x) , f(y) )

Now we prove that h is a bijection .

one - to - one :

Let h(x , y ) = h ( a , b)

\Rightarrow ( f(x) , g(y)) = ( f(a) , g(b))

\Rightarrow f (x) = f(a) and g(y) = g(b)

\Rightarrow x = a and y = b , since f , g are one to one

\Rightarrow ( x , y ) = ( a , b)

So h is one - to-one .

Onto : Let (a , b) \in C ×D

\Rightarrowa \in C , b \in D

As f and g are onto there exist x\in A , y \in B such that

f(x) = a and g(x) = b

\Rightarrow h (x , y ) = ( f(x) , g(y)) = ( a , b)

\Rightarrow ( a , b) has a prem age under h .

So h is onto .

And consequently h is bijection .

Hence ,  A \times B \approx C \times D .

.

.

.

Please comment if needed.